# Learning Latent Transmission and Glare Maps for Lens Veiling Glare Removal

### v1

@@ Abstract
Beyond the commonly recognized optical aberrations, the imaging performance of compact optical systems—including single-lens and metalens designs—is often further degraded by veiling glare caused by stray-light scattering from non-ideal optical surfaces and coatings, particularly in complex real-world environments. This compound degradation undermines traditional lens aberration correction yet remains underexplored. A major challenge is that conventional scattering models (e.g., for dehazing) fail to fit veiling glare due to its spatial-varying and depth-independent nature. Consequently, paired high-quality data are difficult to prepare via simulation, hindering application of data-driven veiling glare removal models. To this end, we propose VeilGen, a generative model that learns to simulate veiling glare by estimating its underlying optical transmission and glare maps in an unsupervised manner from target images, regularized by Stable Diffusion (SD)-based priors. VeilGen enables paired dataset generation with realistic compound degradation of optical aberrations and veiling glare, while also providing the estimated latent optical transmission and glare maps to guide the veiling glare removal process. We further introduce DeVeiler, a restoration network trained with a reversibility constraint, which utilizes the predicted latent maps to guide an inverse process of the learned scattering model. Extensive experiments on challenging compact optical systems demonstrate that our approach delivers superior restoration quality and physical fidelity compared with existing methods. These suggest that VeilGen reliably synthesizes realistic veiling glare, and its learned latent maps effectively guide the restoration process in DeVeiler. All code and datasets will be publicly released at https://github.com/XiaolongQian/DeVeiler.
Figure 1: Compact optical systems suffer from residual aberrations and veiling glare (a), caused by design-induced blur and stray-light scattering from non-ideal surfaces and coatings. (b) Existing methods fail under this compound degradation: a Computational Aberration Correction (CAC) model [liang2021swinir] retrained on aberration-only data fails on unseen veiling glare, while a cascaded state-of-the-art dehazing model [wang2025learning] introduces inconsistent artifacts. Our DeVeiler restores a clean image by jointly correcting the compound degradations.
@@ 1 Introduction
Driven by the demand for compact and high-performance imaging systems in applications like Augmented Reality (AR/VR) [li2021meta, chen2024hybrid] and mobile photography [chen2021extreme_quality, chen_mobile_2023], compact optical systems such as single-lens [qian2025towards] and metalens [tseng2021neural, zhang2025degradation] are becoming increasingly important. However, miniaturized optics inevitably introduce aberrations due to design trade-offs, while low-cost and structural constraints cause surface imperfections and non-ideal coatings, resulting in diffuse stray-light scattering manifested as veiling glare (see Fig. 1(a)). The resulting veiling glare manifests as a widespread and diffuse veil that reduces image contrast even under normal lighting conditions, in contrast to structured artifacts such as lens flare or ghosting, which typically occur when bright sources are in or near the field of view [kotp2023toward_flare_free]. Consequently, residual aberration and veiling glare frequently coexist, forming a compound degradation that severely limits image quality in compact systems. While Computational Aberration Correction (CAC) [heide2013high, chen2021extreme_quality, chen2021optical, li2021universal, jiang2024minimalist, jiang2024flexible, gong2024physics, chen2025physics, li2025optical_attention] can mitigate spatial-varying aberrations, it cannot recover the contrast loss induced by diffuse scattering. Moreover, hardware suppression of veiling glare remains difficult, as compact designs leave little room for baffles or advanced coatings. These factors collectively highlight the urgent need for a computational framework capable of modeling and restoring this coupled optical degradation.
@@
Addressing this compound degradation is non-trivial, as the absence of realistic training data remains a key bottleneck. Physically-accurate simulation of veiling glare demands both complete optomechanical models and computationally prohibitive non-sequential ray-tracing, making large-scale data generation impractical [raskar2008glare, hullin2011physically]. This data scarcity constrains existing paradigms. Classical approaches to veiling glare removal, such as deconvolution-based methods [faulkner1989veiling, talvala2007veiling] or image decomposition [zhang2018single], often rely on restrictive assumptions like spatial invariance and tend to fail in low-texture scenes. Modern learning-based methods that rely on simplified 2D synthesis models [shoshin2021veiling] suffer from a significant domain gap. Moreover, methods designed for visually similar degradations are fundamentally incompatible. Flare removal algorithms [dai2022flare7k, dai2024flare7k++, tsai2025lightsout, FlareX_lishenqu] are ineffective as they target structured artifacts, not the diffuse nature of veiling glare. Dehazing methods [he2010single, wu2023ridcp, wang2024ucl, wang2025learning] are similarly unsuitable, as their atmospheric scattering models are incompatible with the depth-independent physics of glare originating within the lens assembly.
@@
To address this data challenge, recent Stable Diffusion (SD)-based generative models [yang2023synthesizing, peng2024towards, wang2025learning] have been explored, yet they often operate as “black-boxes”, lacking physical grounding. In contrast, we introduce VeilGen, a physics-informed SD-based model that embeds the formation principles of veiling glare directly into the generative process (Fig. 2). In our setting, paired data are only available for aberration correction, while unpaired degraded images from the compact systems exhibit veiling glare on top of residual aberrations, with physical components that are difficult to measure precisely. To bridge this gap, VeilGen incorporates a Latent Optical Transmission and Glare Map Predictor (LOTGMP) that estimates the two key latent components, the transmission and glare maps, within the diffusion process. The latent maps are injected via the Veiling Glare Imposition Module (VGIM) to modulate features, enabling VeilGen to synthesize realistic paired data (see §3.3).
@@
Building upon our generative foundation, we propose DeVeiler, a restoration network trained with a reversibility constraint (Fig. 3(b)). Unlike pure blind-learning approaches, which produce an averaged, suboptimal removal for spatial-varying veiling glare, DeVeiler is structurally guided to learn an inverse mapping of the degradation. However, directly using VeilGen during restoration training is impractical because its multi-step diffusion sampling is computationally expensive. We therefore distill the generative process into a lightweight forward model that preserves its behavior and latent-map conditioning, and use it to supervise the training process of DeVeiler (see §3.4). Its core component, the Veiling Glare Compensation Module (VGCM), uses internally estimated latent maps to modulate image features, thereby precisely reversing the forward glare imposition process (see §3.5). This allows DeVeiler to remove glare based on its underlying physical causes instead of spurious statistical correlations, achieving superior restoration results, as shown in Fig. 1(b). We verify our framework on two challenging optical systems: a large-aperture single-lens and a meta-refractive hybrid lens, both of which exhibit significant compound degradations. Extensive experimental results show that our approach achieves state-of-the-art results, outperforming existing methods. This validates our framework, where a learned forward scattering model provides powerful supervision to ensure its inverse restoration network can effectively utilize the shared latent maps to remove veiling glare. In a nutshell, our contributions are summarized as follows:
@@
Figure 2: Overall architecture of the proposed VeilGen. In Stage I, VeilGen is trained to synthesize compound degradations by using a Latent Optical Transmission and Glare Map Predictor (LOTGMP) to estimate latent maps. These maps then guide the diffusion process via the Veiling Glare Imposition Module (VGIM). $Z_{de}^{\mathcal{T}}$ denotes the target degraded latent representation. $Z_{t}$ denotes the noisy latent representation at timestep $t$ of the forward diffusion process. $Z_{null}$ represents an all-zero latent representation. $txt^{\mathcal{S}/\mathcal{T}}$ denotes the text prompts for the source and target domains, respectively.
@@ 2 Related Work
Computational Aberration Correction. Computational Aberration Correction (CAC) employs algorithms to compensate residual optical aberrations. Recent learning-based approaches [peng2019learned, chen2021extreme_quality, chen2021optical, jiang2024minimalist, luo2024correcting, zhang2025degradation, mandal2025enabling] have surpassed traditional optimization-based methods [wiener1949extrapolation, fish1995blind, schuler2012blind, 2015blind] in performance. However, their dependence on simulated training data leads to a persistent sim-to-real gap and limited generalization across lenses. Efforts toward universal CAC frameworks [li2021universal, gong2024physics, jiang2024flexible] improve cross-lens adaptability by leveraging diverse lens libraries, yet they primarily target aberration-induced blur. The concurrent veiling glare from diffuse scattering remains underexplored, constraining performance under compound degradations.
@@
Veiling Glare Removal. Research on glare-related artifacts generally falls into two categories. Lens flare removal [wu2021how_to_train, dai2022flare7k, dai2023nighttime, dai2024flare7k++, vasluianu2024sfnet, tsai2025lightsout, zhu2025pbfg, FlareX_lishenqu], targets structured, localized artifacts such as bright spots and streaks. In contrast, veiling glare manifests as a diffuse degradation that reduces image contrast, making flare-oriented approaches ineffective. Prior work relied on deconvolution using the Glare Spread Function (GSF) [talvala2007veiling] or on image decomposition [zhang2018image], yet these methods impose restrictive assumptions, often require multiple captures, or fail in low texture scenes. More recent learning-based techniques are limited by the lack of realistic paired data, as existing synthesis strategies [shoshin2021veiling] remain overly simplified. Moreover, previous studies typically handle veiling glare and aberration as independent degradations, whereas our framework jointly addresses both within a unified computational model.
@@
Domain Adaptation in Low-level Vision. Domain Adaptation (DA) is key to mitigating domain shifts in low-level vision tasks [wei2021unsupervised, wang2021unsupervised, li2019semi, chen2021psd], yet generic approaches struggle when the shift arises from compound degradations that couple optical aberrations and veiling glare. This has motivated specialized DA methods for CAC that adapt pre-trained models to new lenses [jiang2025qdmr, jiang2024flexible]. A promising direction is target-domain data synthesis via generative models [yang2023synthesizing, peng2024towards, wang2025learning], with SD-based frameworks such as HazeGen [wang2025learning] showing strong potential. However, for optical degradations with well-defined physical properties, a black-box approach may be suboptimal. To this end, our work introduces a physics-informed generative model that explicitly estimates latent maps for veiling glare. These maps serve dual roles: conditioning an SD-based generator to produce realistic degradations and guiding the restoration network to better remove veiling glare.
@@ 3.1 Problem Formulation
This work aims to restore a latent clean image $I_{c}$ from a single degraded observation $I_{de}$ captured by a compact optical system. The degradations central to this work are twofold: (i) spatial-varying aberration blur and (ii) veiling glare. We model the forward degradation process in two sequential steps. Following the widely adopted patch-wise convolution model [li2021universal, yang2023aberration, luo2024correcting, yang2025efficient], the aberration image patch $I_{ab}^{p}$ is obtained by convolving each clean patch $I_{c}^{p}$ with its corresponding Point Spread Function (PSF) $K^{p}$, representing local aberration characteristics. This aberration patch $I_{ab}^{p}$ is then sequentially degraded by veiling glare according to scattering model, formulated as a combination of attenuation and additive light [zhang2018single, shoshin2021veiling]: each patch $I_{ab}^{p}$ is modulated by a transmission map $T^{p}$ and combined with a glare map $I_{g}^{p}$, where $T^{p}$ denotes local contrast attenuation and $I_{g}^{p}$ represents the veiling glare. For each color channel (omitted for brevity), the overall degradation model is expressed as:
@@ 3.3 Stage I: Realistic Degradation Generation
To overcome data scarcity, we introduce VeilGen, an SD-based generative model that synthesizes realistic paired data (Fig. 2(a)). Built upon DiffBIR [lin2024diffbir] with a fixed Stable Diffusion backbone and IRControlNet, VeilGen adapts the standard denoising paradigm [ho2020denoising] and incorporates a novel mechanism for modeling veiling glare.
Physics-informed Hybrid Training. To bridge a source domain $\mathcal{S}$ with paired, glare-free data and a target domain $\mathcal{T}$ with unpaired compound degradations, we propose a physics-informed hybrid training paradigm inspired by recent work in domain adaptation [wang2025learning]. Since the spatial-varying transmission and glare maps governing veiling glare are difficult to obtain directly, the diffusion model alone cannot capture the underlying physics. To address this, we introduce the Latent Optical Transmission and Glare Map Predictor (LOTGMP), which infers these latent maps during denoising, providing physically meaningful guidance to the generative process. Unlike conventional end-to-end degradation learning, VeilGen steers generation using unified latent maps $c_{vg}=(z_{trans},z_{glare})$:
@@
For the source domain, we use fixed maps ($\mathbf{1},\mathbf{0}$) corresponding to an identity transformation (no veiling glare). For the target domain, we introduce LOTGMP, a lightweight, time-aware network that predicts the latent maps from a noisy latent $z_{t}$, a target degraded latent $z_{de}^{\mathcal{T}}$, and the timestep $t$. The predicted maps are injected via VGIM (Fig. 2(b)), mirroring the forward degradation model (Eq. 1). The overall training objective is a weighted combination of losses from both domains, each conditioned on a domain-specific text prompt:
@@
Degradation-Clean Paired Data Generation. Once trained, VeilGen is employed to generate the paired dataset for training our restoration network in an offline process (Algorithm 1). We use the standard notation from DDPM [ho2020denoising], where $\beta_{t}\in(0,1)$ is the variance schedule, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ is the cumulative product. To generate each degraded image $I_{de}$, we first extract latent maps $c_{vg}$ from an unpaired target-domain image $I_{de}^{\mathcal{T}}$ via the frozen LOTGMP. These maps then guide a full diffusion sampling process that applies the degradation to a clean image $I_{c}$. This sampling process incorporates the blended strategy from [wang2025learning], mixing conditional and unconditional predictions to enhance realism and diversity. This process produces a high-fidelity paired dataset in which each degraded image contains realistic compound degradations, including aberration and veiling glare.

### v2

@@ Abstract
Beyond the commonly recognized optical aberrations, the imaging performance of simplified optical systems—including single-lens and metalens designs—is often further degraded by veiling glare caused by stray-light scattering from non-ideal optical surfaces and coatings, particularly in complex real-world environments. This compound degradation undermines traditional lens aberration correction yet remains underexplored. A major challenge is that conventional scattering models (e.g., for dehazing) fail to fit veiling glare due to its spatial-varying and depth-independent nature. Consequently, paired high-quality data are difficult to prepare via simulation, hindering application of data-driven veiling glare removal models. To this end, we propose VeilGen, a generative model that learns to simulate veiling glare by estimating its underlying optical transmission and glare maps in an unsupervised manner from target images, regularized by Stable Diffusion (SD)-based priors. VeilGen enables paired dataset generation with realistic compound degradation of optical aberrations and veiling glare, while also providing the estimated latent optical transmission and glare maps to guide the veiling glare removal process. We further introduce DeVeiler, a restoration network trained with a reversibility constraint, which utilizes the predicted latent maps to guide an inverse process of the learned scattering model. Extensive experiments on challenging simplified optical systems demonstrate that our approach delivers superior restoration quality and physical fidelity compared with existing methods. These suggest that VeilGen reliably synthesizes realistic veiling glare, and its learned latent maps effectively guide the restoration process in DeVeiler. All code and datasets will be publicly released at https://github.com/XiaolongQian/DeVeiler.
Figure 1: Simplified optical systems suffer from residual aberrations and veiling glare (a), caused by design-induced blur and stray-light scattering from non-ideal surfaces and coatings. (b) Existing methods fail under this compound degradation: a Computational Aberration Correction (CAC) model [29] retrained on aberration-only data fails on unseen veiling glare, while a cascaded state-of-the-art dehazing model [49] introduces inconsistent artifacts. Our DeVeiler restores a clean image by jointly correcting the compound degradations.
@@ 1 Introduction
Driven by the demand for compact and high-performance imaging systems in applications like Augmented Reality (AR/VR) [27, 3] and mobile photography [4, 6], simplified optical systems such as single-lens [38] and metalens [46, 62] are becoming increasingly important. However, simplified optics inevitably introduce aberrations due to design trade-offs, while low-cost and structural constraints cause surface imperfections and non-ideal coatings, resulting in diffuse stray-light scattering manifested as veiling glare (see Fig. 1(a)). The resulting veiling glare manifests as a widespread and diffuse veil that reduces image contrast even under normal lighting conditions, in contrast to structured artifacts such as lens flare or ghosting, which typically occur when bright sources are in or near the field of view [24]. Consequently, residual aberration and veiling glare frequently coexist, forming a compound degradation that severely limits image quality in simplified systems. While Computational Aberration Correction (CAC) [16, 4, 5, 26, 20, 21, 14, 2, 28] can mitigate spatial-varying aberrations, it cannot recover the contrast loss induced by diffuse scattering. Moreover, hardware suppression is highly constrained: applying advanced coatings to such lenses is manufacturing-prohibitive, and baffles fail against in-field stray light. These factors highlight the urgent need for a computational framework capable of modeling and restoring this coupled optical degradation.
@@
Addressing this compound degradation is non-trivial, as the absence of realistic training data remains a key bottleneck. Physically-accurate simulation of veiling glare demands both complete optomechanical models and computationally prohibitive non-sequential ray-tracing, making large-scale data generation impractical [39, 18]. This data scarcity constrains existing paradigms. Classical approaches to veiling glare removal, such as deconvolution-based methods [12, 43] or image decomposition [64], often rely on restrictive assumptions like spatial invariance and tend to fail in low-texture scenes. Modern learning-based methods that rely on simplified 2D synthesis models [42] suffer from a significant domain gap. Moreover, methods designed for visually similar degradations are fundamentally incompatible. Flare removal algorithms [8, 10, 45, 31] are ineffective as they target structured artifacts, not the diffuse nature of veiling glare. Dehazing methods [15, 56, 51, 49] are similarly unsuitable, as their atmospheric scattering models are incompatible with the depth-independent physics of glare originating within the lens assembly.
@@
To address this data challenge, recent Stable Diffusion (SD)-based generative models [58, 36, 49] have been explored, yet they often operate as “black-boxes”, lacking physical grounding. In contrast, we introduce VeilGen, a physics-informed SD-based model that embeds glare formation principles directly into the generative process (Fig. 2). In our setting, paired data are only available for aberration correction, while unpaired degraded images from the simplified systems exhibit veiling glare on top of residual aberrations, with physical components that are difficult to measure precisely. To bridge this gap, VeilGen incorporates a Latent Optical Transmission and Glare Map Predictor (LOTGMP) that estimates the two key latent components, the transmission and glare maps, within the diffusion process. The latent maps are injected via the Veiling Glare Imposition Module (VGIM) to modulate features, enabling VeilGen to synthesize realistic paired data (see §3.3).
@@
Building upon our generative foundation, we propose DeVeiler, a restoration network trained with a reversibility constraint (Fig. 3(b)). Unlike pure blind-learning approaches, which produce an averaged, suboptimal removal for spatial-varying veiling glare, DeVeiler is structurally guided to learn an inverse mapping of the degradation. However, directly using VeilGen during restoration training is impractical because its multi-step diffusion sampling is computationally expensive. We therefore distill the generative process into a lightweight forward model that preserves its behavior and latent-map conditioning, and use it to supervise the training process of DeVeiler (see §3.4). Its core component, the Veiling Glare Compensation Module (VGCM), uses internally estimated latent maps to modulate image features, thereby precisely reversing the forward glare imposition process (see §3.5). This allows DeVeiler to remove glare based on its underlying physical causes instead of spurious statistical correlations, achieving superior restoration results, as shown in Fig. 1(b). While our ultimate goal is addressing the compound degradation of simplified optics, our core technical novelty focuses specifically on veiling glare removal. Since aberrations are sufficiently mitigated via foundational source-domain supervision, the under-constrained veiling glare remains the critical bottleneck (detailed in the Supplement). Validated on two challenging lens prototypes, our framework achieves state-of-the-art restoration by effectively utilizing shared latent maps to explicitly model and reverse this physical glare. In a nutshell, our contributions are summarized as follows:
@@ 2 Related Work
Figure 2: Overall architecture of the proposed VeilGen. In Stage I, VeilGen is trained to synthesize compound degradations by using a Latent Optical Transmission and Glare Map Predictor (LOTGMP) to estimate latent maps. These maps then guide the diffusion process via the Veiling Glare Imposition Module (VGIM). $Z_{de}^{\mathcal{T}}$ is obtained by encoding target degraded images via the frozen VAE. $Z_{t}$ denotes the noisy latent representation at timestep $t$ of the forward diffusion process. $Z_{null}$ represents an all-zero latent representation. $txt^{\mathcal{S}/\mathcal{T}}$ denotes the text prompts for the source and target domains, respectively.
@@
Computational Aberration Correction. Computational Aberration Correction (CAC) employs algorithms to compensate residual optical aberrations. Recent learning-based approaches [37, 4, 5, 20, 33, 62, 34] now outperform traditional optimization-based methods [54, 13, 41, 61]. However, their reliance on simulated data causes a persistent sim-to-real gap and limits generalization across lenses. Efforts toward universal CAC frameworks [26, 14, 21] improve cross-lens adaptability using diverse lens libraries, yet they primarily target aberration-induced blur. The concurrent veiling glare remains underexplored, constraining performance under compound degradations.
@@
Veiling Glare Removal. Research on glare-related artifacts generally falls into two categories. Lens flare removal [57, 8, 9, 10, 47, 45, 65, 31], targets structured, localized artifacts such as bright spots and streaks. In contrast, veiling glare manifests as a diffuse degradation that reduces image contrast, making flare-oriented approaches ineffective. Prior work relied on deconvolution using the Glare Spread Function (GSF) [43] or on image decomposition [64], yet these methods impose restrictive assumptions, often require multiple captures, or fail in low texture scenes. More recent learning-based techniques are limited by the lack of realistic paired data, as existing synthesis strategies [42] remain overly simplified. Moreover, previous studies typically handle veiling glare and aberration as independent degradations, whereas our framework jointly addresses both within a unified computational model.
@@
Domain Adaptation in Low-level Vision. While Domain Adaptation (DA) mitigates domain shifts in low-level vision [53, 50, 25, 7], generic approaches struggle with compound degradations coupling optical aberrations and veiling glare. This motivates specialized DA methods adapting pre-trained models to novel lenses [22, 21]. Although target-domain synthesis via SD-based generative models [58, 36, 49] shows promise, black-box approaches remain suboptimal for well-defined physical optical degradations. To this end, our work introduces a physics-informed generative model that explicitly estimates latent maps for veiling glare. These maps serve dual roles: conditioning an SD-based generator to produce realistic degradations and guiding the restoration network to better remove veiling glare.
@@ 3.1 Problem Formulation
This work aims to restore a latent clean image $I_{c}$ from a degraded observation $I_{de}$ captured by a simplified optical system. The degradations central to this work are twofold: (i) spatial-varying aberration blur and (ii) veiling glare. We model the forward degradation process in two sequential steps. Following the widely adopted patch-wise convolution model [26, 59, 33, 60], the aberration image patch $I_{ab}^{p}$ is obtained by convolving each clean patch $I_{c}^{p}$ with its corresponding Point Spread Function (PSF) $K^{p}$, representing local aberration characteristics. This aberration patch $I_{ab}^{p}$ is then sequentially degraded by veiling glare according to scattering model, formulated as a combination of attenuation and additive light [64, 42]: each patch $I_{ab}^{p}$ is modulated by a transmission map $T^{p}$ and combined with a glare map $I_{g}^{p}$, where $T^{p}$ denotes local contrast attenuation and $I_{g}^{p}$ represents the veiling glare. For each color channel (omitted for brevity), the overall degradation model is expressed as:
@@ 3.3 Stage I: Realistic Degradation Generation
To overcome data scarcity, we introduce VeilGen, an SD-based generative model that synthesizes realistic paired data (Fig. 2(a)). Built upon DiffBIR [30] with a fixed Stable Diffusion backbone and IRControlNet, VeilGen adapts the standard denoising paradigm [17] and incorporates a novel mechanism for modeling veiling glare.
Physics-informed Hybrid Training. To bridge a source domain $\mathcal{S}$ with paired, glare-free data and a target domain $\mathcal{T}$ with unpaired compound degradations, we propose a physics-informed hybrid training paradigm inspired by recent work in domain adaptation [49]. Since the spatial-varying transmission and glare maps governing veiling glare are difficult to obtain directly, the diffusion model alone cannot capture the underlying physics. To address this, we introduce the Latent Optical Transmission and Glare Map Predictor (LOTGMP), which infers these latent maps during denoising, providing physically meaningful guidance to the generative process. Unlike conventional end-to-end degradation learning, VeilGen steers generation using unified latent maps $c_{vg}=(z_{trans},z_{glare})$:
@@
For the source domain, we use fixed maps ($\mathbf{1},\mathbf{0}$) corresponding to an identity transformation (no veiling glare). For the target domain, the proposed LOTGMP predicts the latent maps from a noisy latent $z_{t}$, a target degraded latent $z_{de}^{\mathcal{T}}$, and the timestep $t$. The predicted maps are injected via VGIM (Fig. 2(b)), mirroring the forward degradation model (Eq. 1). The overall training objective is a weighted combination of losses from both domains, each conditioned on a domain-specific text prompt:
@@
Degradation-Clean Paired Data Generation. Once trained, VeilGen is employed to generate the paired dataset for training our restoration network in an offline process (Algorithm 1). We use the standard notation from DDPM [17], where $\beta_{t}\in(0,1)$ is the variance schedule, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ is the cumulative product. To generate each degraded image $I_{de}$, we first extract latent maps $c_{vg}$ from an unpaired target-domain image $I_{de}^{\mathcal{T}}$ via the frozen LOTGMP. These maps then guide a full diffusion sampling process that applies the degradation to a clean image $I_{c}$. This sampling process incorporates the blended strategy from [49], mixing conditional and unconditional predictions to enhance realism and diversity. This process produces a high-fidelity paired dataset in which each degraded image contains realistic compound degradations, including aberration and veiling glare.